OpenStudy (anonymous):

Trigonometry Help (only 4 left!) Prove the Identity: sec^2 x= (4sin^2 x)/(sin^2 x)

OpenStudy (anonymous):

are you sure the question is complete and correct

OpenStudy (anonymous):

 is what the problem looks like

OpenStudy (anonymous):

now the question is different from what you posted

OpenStudy (anonymous):

yeah a tiny bit, go by the picture, the picture is correct. that was an accident

OpenStudy (anonymous):

do u know sin2x = 2sinx cosx

OpenStudy (anonymous):

sin 2x or sin^2 x?

OpenStudy (anonymous):

nvm sin 2x

OpenStudy (anonymous):

so sec^2 x = 1/cos^2 x = (1/cos^2 x) * (sin^2 x/sin^2 x )

OpenStudy (anonymous):

(sin^2 x/( cos^2 x *sin^2 x )) =(4*sin^2 x/(4* cos^2 x *sin^2 x )) = 4*sin^2 x /(2sinx cosx)^2 = 4*sin^2 x /(sin2x)^2

OpenStudy (anonymous):

hope it is clear to you